OpenStudy (xmoses1):

The y-intercept of the line whose equation is 2x - 3y = 6 is -2 3 6

OpenStudy (xmoses1):

@ganibl

OpenStudy (xmoses1):

and my last question i have is: The slope of the line whose equation is 5x - 3y = 4 is 5/3 4/5 -3/5

OpenStudy (anonymous):

Lets first convert this into slope intercept form that is of the form y = mx + b

OpenStudy (anonymous):

I will do the first one and u try the second one in the same way is that fine?

OpenStudy (xmoses1):

so y=2/3x-6

OpenStudy (xmoses1):

yeah i think i can do it :)

OpenStudy (anonymous):

Thats right :)

OpenStudy (anonymous):

So now u got the answer for first one :)

OpenStudy (xmoses1):

ok and then same process for the second one?

OpenStudy (anonymous):

Yes exactly :)

OpenStudy (xmoses1):

hold on my answer choices for the first Q are -2 3 6, did i forget a step?

OpenStudy (anonymous):

No U didnt u got there :) Any equation of the form y = mx + b the value of b is called as y-intercept

OpenStudy (xmoses1):

oh oops lol

OpenStudy (anonymous):

so now in ur first one u got the equation y = 2/3x - 6 with b value as -6 which is ur y-intercept

OpenStudy (xmoses1):

but i would have to make it a positive 6 though

OpenStudy (anonymous):

No it is negative 6 only Its not mandatory for u to make it positive

OpenStudy (xmoses1):

oh ok thank you:)

OpenStudy (anonymous):

y = mx + b is the standard form and b can be either negative or positive

OpenStudy (anonymous):

yw :)